# Physical Human-Robot Interaction: A Critical Review of Safety Constraints

### v2

@@ 1 Introduction
To provide practical guidance, the International Organization for Standardization (ISO) helps to bridge the gap between the abstract concept of safety and its operational implementation by adopting a functional perspective in which safety is regarded as the capability of a robotic system to operate within limits that ensure risks to humans are reduced to an acceptable level. These limits are typically expressed in terms of forces, velocities, accelerations, energy transfer, or spatial constraints, thereby providing quantifiable and enforceable metrics for safe operation. In 2011, the ISO released the standard ISO 10218-2 [25], which established foundational safety requirements for the integration of industrial robots into human work environments. Building on this, the technical specification ISO/TS 15066 [26] was introduced in 2016 to support practitioners in ensuring safety in HRI. A concrete risk comes from the belief that cobots’ design makes them inherently safe, resulting in inadequate control measures, especially in dynamic and complex industrial settings. To this purpose, in 2025, ISO 10218-2 was revised to deleverage on the concept of cobots in general: only the application can be designed, verified, and validated as collaborative, rather than the robot [27]. Besides this conceptual shift, for what concerns the safety aspect, these new specifications reaffirm the core practice introduced in ISO/TS 15066, which is still the main technical guideline in the field. The technical specification ISO/TS 15066 provides quantitative limits to mitigate the risk of human injury during collaborative operations. It enables the dynamic coexistence of humans and robots in the same workspace by defining the allowable motion characteristics under four different collaborative operation modes. Among these modes, the only one that accounts for physical interaction between the robot and a human is the Power and Force Limiting (PFL) mode, which presents a pre-collision guideline aimed at producing safe pHRI.
@@
In this perspective, the present paper aims to provide a clearer and more rigorous understanding of commonly recognized safety constraints in pHRI, focusing on how they are derived and the assumptions underlying their formulation. Understanding the core rationale behind the results and making these assumptions explicit is essential for correctly interpreting the standards, applying them coherently, and addressing ambiguous or underspecified situations with a critical and informed view. This detailed examination also provides the foundation for analyzing the constraints’ impact on both safety guaranties and robot performance, ultimately guiding the identification of informed improvements and best practices.
@@ 2 Operational Constraints of Robotic Motion
A fundamental preliminary step consists of establishing a model for human-robot contact. In accordance with ISO/TS 15066 and a substantial portion of the relevant literature, the interaction is typically idealized as a collision between two lumped masses interconnected by a compliant element, as depicted in 2(a). The robot dynamics are reduced to an effective mass $m_{R}$, which denotes the configuration-dependent reflected inertia of the manipulator at the point of contact. The impacted region of the human body is represented by an effective mass $m_{H}$, which characterizes the inertial properties of the corresponding anatomical segment. The deformation of biological tissues at the contact interface is modeled by a linear elastic spring with stiffness $k$, which accounts for the elastic energy stored during the impact event.
@@ 3.1 Derivation process and fully inelastic assumptions
It is important to notice that the results detailed in Sec. 2 follow directly from the constraint formulation in ISO/TS 15066, while making explicit the intermediate steps required to derive them. We also emphasize that the constraints derived in Sec. 2 define operative limits on the robot’s motion in terms of the maximum admissible speed (or, equivalently, kinetic energy) before impact , such that the force or pressure generated during the collision does not exceed the tolerable pain thresholds for the contacted human body region. By representing the interaction as a collision between two discrete masses and explicitly accounting for the local mechanical stiffness of the impacted human body region, the critical time instant for safety assessment corresponds to the moment of maximum deformation, at which the biomechanical loading reaches its peak value. The entire derivation pipeline, shown in Figure 3, is formulated at this instant. As a result, the only phase relevant for safety is the time interval before maximum compression. The resulting constraints do not require nor depend on any assumptions about what happens after this peak load. Unsafe modes are triggered by the maximum deformation alone; whether the collision dissipates energy (inelastic) or returns it (elastic) is irrelevant because such post-maximum-load dynamics occur only after the critical threshold has already been reached. In contrast, ISO/TS 15066 and prior works such as [49, 35] derive the limits from the difference between the kinetic energy before and after impact, explicitly assuming a fully inelastic collision in which all lost kinetic energy is dissipated through tissue deformation or fracture. This assumption is strong and not generally representative, as real human-robot impacts occur between the extremes of perfectly elastic and perfectly inelastic behavior, raising concerns about the conservativeness of the fully inelastic model [6]. In summary, the robot’s motion constraints (identical to the one reported in ISO/TS 15066) should be interpreted as an instantaneous safety condition: the robot must enter the collision with sufficiently low kinetic energy to prevent the tissue from exceeding its allowable maximum compression. The restitution characteristics of the impact, regardless of how elastic or inelastic they may be, do not alter this requirement. Our formulation makes this independence explicit and avoids assumptions about post-impact dynamics altogether.
@@ 3.2 Pain Thresholds
In this remark, we observe from (2) that the maximum admissible speed, $v_{0,\max}$, exhibits a linear dependence on the force bound $F_{\max}$ and an inverse dependence on the stiffness parameter $k$. Consequently, if a collision involves a body region characterized by lower effective stiffness or higher maximum tolerable force, the robot can operate at a higher allowable velocity, since increased compliance or more relaxed force limits permit greater energy transfer without surpassing safety thresholds. Furthermore, it should be emphasized that the maximum admissible pressure $p_{\max}$ is typically determined from biomechanical studies, whereas the corresponding maximum allowable interaction force $F_{\max}$ is obtained according to $F_{\max}=Ap_{\max}$ for a specific contact area $A$, which in turn depends on the robot’s geometry, its end-effector equipment, and its specific configuration during physical interaction. Directly using the $F_{\max}$ values reported in technical specifications may appear more convenient, as it removes the need for an explicit determination of $A$, which can indeed be difficult to assess. However, these tabulated values are typically computed under the assumption of a conservative nominal contact area of $A=1\mathrm{cm^{2}}$. Consequently, when a more accurate estimate of the actual contact area is available, it is advantageous to adjust $F_{\max}$ accordingly. It is important to note that the admissible robot velocity exhibits a linear dependence on the contact area. As a result, for instance, doubling the contact area to $2\mathrm{cm^{2}}$ allows the robot to operate at twice the velocity associated with the nominal $F_{\max}$ specified for $A=1\mathrm{cm^{2}}$. Collaborative robots are intentionally designed to minimize sharp edges and protrusions, which reduces local pressure and lowers the risk of injury. However, when additional equipment (e.g., grippers or tools) is attached to the robot, or when sharp objects are being manipulated, the effective contact area can be reduced, thereby lowering the maximum safe force that can be applied during a collision. In such situations, control strategies and trajectory planning can be used to increase safety.
@@ 3.3 Human Effective Mass
In the limiting case, i.e., for $m_{H}=\infty$, this situation can be interpreted as a clamped condition, in which the human body segment is constrained and therefore unable to recoil in response to the impact, as depicted in Figure 4. The reduced-mass term in the velocity expression reaches its minimum possible value, since any finite human mass would contribute an additional $1/m_{H}$ to the sum, thereby increasing $v_{0,\max}$. Therefore, assuming the possibility of a fully clamped human provides the strictest, most conservative upper bound on robot velocity. In this case, the admissible impact velocity depends exclusively on the robot effective mass and the tissue stiffness. A lighter robot reduces the inertia that must be absorbed by the tissue, directly increasing $v_{0,\max}$. Physically, nearly all of the kinetic energy that is transferred from the robot to the human during the collision is stored as elastic energy in the tissues, with negligible motion of the human segment. The corresponding maximum initial kinetic energy of the robot is then approximately equal to the tissue energy capacity. This scenario is thus more conservative than the general case in which the human operator is able to recoil: for the same robot mass and tissue stiffness, allowing the human segment to move ($m_{H}$ finite) increases the reduced-mass term in the velocity formula, which in turn increases $v_{0,\max}$. Consequently, designing the robot to respect $v_{0,\max}$ under the potential-clamp assumption ensures that peak tissue forces are not exceeded under any realistic human mobility condition, providing a robust safety margin.
@@ 4.2 Safe Control Strategies
Whether safety hazard severity thresholds are defined directly in terms of energy, as in Sec. 4.1, or inferred from force or pressure limits by mapping them to an equivalent potential energy, as in Sec. 2, the resulting methods ultimately constrain the robot’s kinetic energy. This has motivated a substantial body of research on energy-centric safety mechanisms for pHRI. Such mechanisms have the significant benefit of being inherently compatible with passivity-based control schemes [58], thereby facilitating formal stability proofs, enabling the coordinated design of task and safety objectives, and allowing stable interaction with unknown or uncertain environments. Within the context of these energy-based control approaches, the notion of passivity plays a central role. Passivity is a system-theoretic property that, roughly speaking, ensures that a dynamical system does not autonomously generate an unbounded amount of energy [58]. This property is stronger than standard notions of stability and is highly desirable for closed-loop control in pHRI for several reasons [55]. A particularly relevant aspect for pHRI is that the environment with which the robot interacts cannot, in general, be assumed to be known or accurately modeled during task execution. By enforcing passivity at the robot level, stability of the closed-loop system can be guaranteed for the interconnection with any passive environment, independently of its specific dynamics.
@@ 5 Final Discussion
Figure 5: Distribution of admissible end-effector velocities across ISO body regions for three safety formulations: transient (T) contact with scaled force/pressure limits (green), quasi-static (QS) contact following ISO/TS 15066 limits (orange), and the potentially-clamped scenario using (6) (blue). Boxplots show the variability of velocity limits resulting from the configuration- and direction-dependent apparent mass $m_{\mathbf{u}}(q)$, while star markers ($\bigstar$) indicate the corresponding scalar limits obtained using the constant ISO-based mass $m_{R}^{ISO}$.
@@
In Figure 5, we report the pre-collision speed limits across ISO-defined body regions under different safety constraint options. In particular, we compare quasi-static contact conditions, which include potentially-clamped scenarios, with transient contact conditions (in which clamping is excluded). For the quasi-static contact, two alternative constraints are considered. The first follows the approach suggested in ISO/TS 15066, which applies the kinetic energy constraint in (4) without explicit assumptions on human mobility, i.e., $\Delta K\leq U_{s,\max}^{QS}$, where $U_{s,\max}^{QS}$ denotes the maximum admissible elastic energy derived from quasi-static thresholds. The second adopts the model in (6), which explicitly assumes an immobilized human body part. In this case, the entire pre-impact kinetic energy may be transferred to the human, leading to the stricter condition $K_{0}\leq U_{s,\max}^{QS}$. For the transient contact, we consider the standard constraint (4) as specified in ISO/TS 15066, i.e., $\Delta K\leq U_{s,\max}^{T}$, where $U_{s,\max}^{T}$ is derived from force and pressure limits scaled with respect to quasi-static values. In accordance with the technical specification, these transient limits are obtained by applying a multiplicative factor of two to the quasi-static values for all body regions, with the exception of the skull/forehead and the face, for which no scaling is applied. Within the same figure, we also compare speed limits obtained using two different estimates of the robot’s effective mass: the constant value $m_{R}^{ISO}$ and the configuration-dependent apparent mass $m_{\mathbf{u}}(q)$. While $m_{R}^{ISO}$ yields a constant speed limit for each body region, $m_{\mathbf{u}}(q)$ varies with the robot’s configuration and contact direction. Consequently, the latter is reported as a statistical distribution, visualized through boxplot representations.
@@
When human motion and environmental context cannot be reliably tracked, the robot must operate under fixed and conservative safety limits. We consider as baseline reference values the average admissible pre-collision speed limits $v_{\mathrm{max}}^{avg}$ obtained under the transient contact assumption (i.e., no clamping expected), computed using the configuration- and direction-dependent robot effective mass $m_{R}=m_{\mathbf{u}}(q)$. Table 1 reports, for each human body region, these baseline speed limits. The table also shows how these limits are modified under more conservative design assumptions: (i) using the constant effective mass approximation from ISO/TS 15066 ($m_{R}=m_{R}^{ISO}$), and (ii) assuming that a clamped condition may occur, enforced through the constraint $K_{0}\leq U_{s,\max}^{QS}$. These values quantify the reduction in achievable speed when either the robot’s inertia is over-approximated or human mobility is neglected. The largest reductions occur for peripheral and highly mobile segments, such as the hands, fingers, and arms, where admissible speeds can fall below $15\%$ of the baseline, whereas larger, less mobile regions like the torso and pelvis are comparatively less affected. The values reported in Table 2 further quantify the additional performance loss that arises when the robot controller lacks information about the specific human body region that may be involved in a collision. In the absence of body-part awareness, safety limits must be enforced according to the most restrictive body region, namely the face. As a consequence, all potential contacts inherit face-based limits, resulting in substantial reductions in allowable speed. Under the ISO effective mass approximation $m_{R}=m_{R}^{ISO}$, this corresponds to maximum speeds of approximately $0.15\,\mathrm{m/s}$ for transient contact using (5), and $0.10\,\mathrm{m/s}$ for quasi-static contact using (6). Table 2 shows that, under worst-case assumptions, admissible speeds drop below $70\%$ of the baseline in transient contact, and below $46\%$ in the quasi-static case. For most of the body regions, the resulting speed is below $10\%$ of what could be safely achieved if body-region awareness were available. These results clearly illustrate the performance penalty induced by conservative, worst-case design and motivate the adoption of perception-driven, context-aware safety strategies capable of adapting limits to the actual interaction conditions.

### v3

@@ 1 Introduction
To provide practical guidance, the International Organization for Standardization (ISO) helps to bridge the gap between the abstract concept of safety and its operational implementation by adopting a functional perspective in which safety is regarded as the capability of a robotic system to operate within limits that ensure risks to humans are reduced to an acceptable level. These limits are typically expressed in terms of forces, velocities, accelerations, energy transfer, or spatial constraints, thereby providing quantifiable and enforceable metrics for safe operation. In 2011, the ISO released the standard ISO 10218-2 [25], which established foundational safety requirements for the integration of industrial robots into human work environments. Building on this, the technical specification ISO/TS 15066 [26] was introduced in 2016 to support practitioners in ensuring safety in HRI. A concrete risk comes from the belief that cobots’ design makes them inherently safe, resulting in inadequate control measures, especially in dynamic and complex industrial settings. To this purpose, in 2025, ISO 10218-2 was revised to deleverage on the concept of cobots in general: only the application can be designed, verified, and validated as collaborative, rather than the robot [27]. Besides this conceptual shift, the new standard confirms the core methods for quasi-static and transient contact in Power and Force Limiting mode originally defined in ISO/TS 15066 by incorporating them unchanged into Annex M of ISO 10218-2:2025. As a result, both documents are technically equivalent regarding these safety limits and may be used interchangeably for the purposes of this work. Nevertheless, we continue to cite ISO/TS 15066, as it remains the predominant reference for these technical guidelines in pHRI and future revision will preserve the same identifier. The technical specification ISO/TS 15066 provides quantitative limits to mitigate the risk of human injury during collaborative operations. It enables the dynamic coexistence of humans and robots in the same workspace by defining the allowable motion characteristics under four different collaborative operation modes. Among these modes, the only one that accounts for physical interaction between the robot and a human is the Power and Force Limiting (PFL) mode, which presents a pre-collision guideline aimed at producing safe pHRI.
@@
In this perspective, the present paper aims to provide a clearer understanding of commonly recognized safety constraints in pHRI. This objective is pursued by presenting a clear, self-contained, and didactic reconstruction of the derivation process, thereby offering a coherent and unified view of the concepts and assumptions underlying their formulation, which are often presented in a fragmented manner in the literature. Such clarification is essential for correctly interpreting the standards, applying them consistently, and critically assessing their implications. It also enables a more transparent analysis of the trade-off between safety guarantees and robotic performance, ultimately supporting the identification of more informed design choices and best practices.
@@ 2 Operational Constraints of Robotic Motion
A fundamental preliminary step consists of establishing a model for human-robot contact. In accordance with ISO/TS 15066 and a substantial portion of the relevant literature, the interaction is typically idealized as a collision between two lumped masses interconnected by a compliant element, as depicted in 2(a). Note that this one-dimensional collision model corresponds to a mapping of the interaction between a multi-articulated robotic system and a spatially distributed human body onto the direction of contact. This mapping leads to a scalar formulation in which the interaction dynamics is governed by equivalent masses along the contact direction, representing the portions of the system that effectively participate in the interaction. Specifically, the robot is reduced to an effective mass $m_{R}$, which denotes the configuration- and direction-dependent reflected inertia of the manipulator at the contact point. Analogously, the impacted region of the human body is modeled by an effective mass $m_{H}$, which characterizes the apparent inertia of the corresponding anatomical segment involved in the interaction. Within the same formulation, the deformation of biological tissues at the contact interface is modeled by a linear elastic spring with stiffness $k$, which accounts for the elastic energy stored during the impact event
@@ 3.1 Derivation process and fully inelastic assumptions
It is important to notice that the results detailed in Sec. 2 follow directly from the constraint formulation in ISO/TS 15066, while making explicit the intermediate steps required to derive them. We also emphasize that the constraints derived in Sec. 2 define operative limits on the robot’s motion in terms of the maximum admissible speed (or, equivalently, kinetic energy) before impact , such that the force or pressure generated during the collision does not exceed the tolerable pain thresholds for the contacted human body region. By representing the interaction as a collision between two discrete masses and explicitly accounting for the local mechanical stiffness of the impacted human body region, the critical time instant for safety assessment corresponds to the moment of maximum deformation, at which the biomechanical loading reaches its peak value. The entire derivation pipeline, shown in Figure 3, is formulated at this instant. As a result, the only phase relevant for safety is the time interval before maximum compression. The resulting constraints do not require nor depend on any assumptions about what happens after this peak load. Unsafe modes are triggered by the maximum deformation alone; whether the collision dissipates energy (inelastic) or returns it (elastic) is irrelevant because such post-maximum-load dynamics occur only after the critical threshold has already been reached. In contrast, ISO/TS 15066 and prior works such as [49, 35] derive the limits from the difference between the kinetic energy before and after impact, explicitly assuming a fully inelastic collision in which all lost kinetic energy is dissipated through tissue deformation or fracture. This assumption is strong and not generally representative, as real human-robot impacts occur between the extremes of perfectly elastic and perfectly inelastic behavior, raising concerns about the conservativeness of the fully inelastic model [6]. In summary, the robot’s motion constraints (identical to the one reported in ISO/TS 15066) should be interpreted as an instantaneous safety condition: the robot must enter the collision with sufficiently low kinetic energy to prevent the tissue from exceeding its allowable maximum compression. The restitution characteristics of the impact, regardless of how elastic or inelastic they may be, do not alter this requirement.
@@ 3.2 Pain Thresholds
In this remark, we observe from (2) that the maximum admissible speed, $v_{0,\max}$, exhibits a linear dependence on the force bound $F_{\max}$ and an inverse dependence on the stiffness parameter $k$. Consequently, if a collision involves a body region characterized by lower effective stiffness or higher maximum tolerable force, the robot can operate at a higher allowable velocity, since increased compliance or more relaxed force limits permit greater energy transfer without surpassing safety thresholds. Furthermore, it should be emphasized that the maximum admissible pressure $p_{\max}$ is typically determined from biomechanical studies, whereas the corresponding maximum allowable interaction force $F_{\max}$ is obtained according to $F_{\max}=Ap_{\max}$ for a specific contact area $A$, which in turn depends on the robot’s geometry, its end-effector equipment, and its specific configuration during physical interaction. Directly using the $F_{\max}$ values reported in technical specifications may appear more convenient, as it removes the need for an explicit determination of $A$, which can indeed be difficult to assess. However, these tabulated values are typically computed under the assumption of a conservative nominal contact area of $A\leq 1\mathrm{cm^{2}}$. Consequently, when a more accurate estimate of the actual contact area is available, it is advantageous to adjust $F_{\max}$ accordingly. It is important to note that the admissible robot velocity exhibits a linear dependence on the contact area. As a result, for instance, doubling the contact area to $2\mathrm{cm^{2}}$ allows the robot to operate at twice the velocity associated with the nominal $F_{\max}$ specified for $A=1\mathrm{cm^{2}}$. Collaborative robots are intentionally designed to minimize sharp edges and protrusions, which reduces local pressure and lowers the risk of injury. However, when additional equipment (e.g., grippers or tools) is attached to the robot, or when sharp objects are being manipulated, the effective contact area can be reduced, thereby lowering the maximum safe force that can be applied during a collision. In such situations, control strategies and trajectory planning can be used to increase safety.
@@ 3.3 Human Effective Mass
The limiting case, i.e., for $m_{H}=\infty$, can be interpreted as a clamped condition in which the human body segment is constrained and therefore unable to recoil in response to the impact, as depicted in Figure 4. In this situation, nearly all of the kinetic energy that is transferred from the robot to the human during the collision is stored as elastic energy in the tissues, with no motion of the human segment. The corresponding maximum allowable initial kinetic energy of the robot is therefore approximately equal to the tissue energy capacity. This scenario is thus more conservative than the general case in which the human operator is able to recoil: for the same robot mass and tissue stiffness, allowing the human segment to move ($m_{H}$ finite) increases the reduced-mass term in the velocity formula, which in turn increases $v_{0,\max}$. Consequently, designing the control to respect safety speed (or energy) constraints under the potential-clamp assumption ensures that peak forces at the tissue do not exceed pain-sensitive thresholds under any realistic human mobility condition, providing a robust safety margin.
@@ 4.2 Safe Control Strategies
Whether safety hazard severity thresholds are defined directly in terms of energy, as in Sec. 4.1, or inferred from force or pressure limits by mapping them to an equivalent potential energy, as in Sec. 2, the resulting methods ultimately constrain the robot’s kinetic energy. This has motivated a substantial body of research on energy-centric safety mechanisms for pHRI. Such mechanisms have the significant benefit of being inherently compatible with passivity-based control schemes [58], thereby facilitating formal stability proofs, enabling the coordinated design of task and safety objectives, and allowing stable interaction with unknown or uncertain environments. Within the context of these energy-based control approaches, the notion of passivity plays a central role. Passivity is a system-theoretic property that, roughly speaking, ensures that a dynamical system does not autonomously generate an unbounded amount of energy [58]. This property is stronger than standard notions of stability and is highly desirable for closed-loop control in pHRI for several reasons [55]. A particularly relevant aspect for pHRI is that the environment with which the robot interacts cannot, in general, be assumed to be known or accurately modeled during task execution. By enforcing passivity at the robot level, stability of the closed-loop system can be guaranteed for the interconnection with any passive environment, independently of its specific dynamics. Multiple approaches have also moved beyond a purely passivity-based formulation by explicitly enforcing safety through bounds on the energy transfer required to satisfy the constraint in (4). Consistent with the modeling abstraction introduced in Sec. 2, these methods rely on the same one-dimensional interaction representation, in which the full dynamics of a multi-articulated robotic system in contact with a human body are mapped onto the direction of contact
@@ 5 Final Discussion
Figure 5: Admissible end-effector velocities across ISO body regions for three safety formulations: transient (T) contact with scaled force/pressure limits (green), quasi-static (QS) contact following ISO/TS 15066 limits (orange), and the potentially-clamped scenario using (6) (blue). Boxplots show the variability of velocity limits resulting from the configuration- and direction-dependent apparent mass $m_{\mathbf{u}}(q)$, while star markers ($\bigstar$) indicate the corresponding scalar limits obtained using the constant ISO-based mass $m_{R}^{ISO}$.
@@
In Figure 5, we report the pre-collision speed limits across ISO-defined body regions under different safety constraint options. In particular, we compare quasi-static contact conditions, which include potentially-clamped scenarios, with transient contact conditions (in which clamping is excluded). For the quasi-static contact, two alternative constraints are considered. The first follows the approach suggested in ISO/TS 15066, which applies the kinetic energy constraint in (4) without explicit assumptions on human mobility, i.e., $\Delta K\leq U_{s,\max}^{QS}$, where $U_{s,\max}^{QS}$ denotes the maximum admissible elastic energy derived from quasi-static thresholds. The second adopts the model in (6), which explicitly assumes an immobilized human body part. In this case, the entire pre-impact kinetic energy may be transferred to the human, leading to the stricter condition $K_{0}\leq U_{s,\max}^{QS}$. For the transient contact, we consider the standard constraint (4) as specified in ISO/TS 15066, i.e., $\Delta K\leq U_{s,\max}^{T}$, where $U_{s,\max}^{T}$ is derived from force and pressure limits scaled with respect to quasi-static values. In accordance with the technical specification, these transient limits are obtained by applying a multiplicative factor of two to the quasi-static values for all body regions, with the exception of the skull/forehead and the face, for which no scaling is applied. Within the same figure, we also compare speed limits obtained using two different estimates of the robot’s effective mass: the constant value $m_{R}^{ISO}$ and the configuration-dependent apparent mass $m_{\mathbf{u}}(q)$. While $m_{R}^{ISO}$ yields a constant speed limit for each body region, $m_{\mathbf{u}}(q)$ varies with the robot’s configuration and contact direction. Consequently, the latter is visualized using boxplots, which summarize the spread across the sampled configurations.
@@
[…]